OpenStudy (anonymous):

If 4x^2 + kx + 25 = 0 has exactly one solution, what could be the value of k?

OpenStudy (anonymous):

Quadratic formula: \[\sqrt{b^2-4ac}\] has to equal 0 b^2=4ac b^2=400 b=20

OpenStudy (anonymous):

thanks, but how do you find what a and c are?

OpenStudy (anonymous):

wait i figured it out! Thanks so much!!